I need to your help please solve these problems


1)A total of Rs 14000 invested for a year , part at 5% and the rest at 6.5% . If Rs 760 is the total interest, how much is invested at 5%.


2)Mr javed invested a total of Rs 10,000 in three different accouts , the accounts paid simple interest at an annual rate of 8%, 9%,7.5% respectively. Total interest earned for the year was Rs 845 . The amount in the 9% account was the twice the amount invested in the 7.5% account. How much did Mr.Javed invest in each account.

Answer by stanbon(75887)

1)A total of Rs 14000 invested for a year , part at 5% and the rest at 6.5% . If Rs 760 is the total interest, how much is invested at 5%.
---------
Interest Equation:
0.065x + 0.5(14000-x) = 760
65E + 50*14000 - 50x = 760000
15x = 6000
x = 4000 (amt invested at 6.5%)
14000-x = 10000 (amt. invested at 5%)
----------------------------------------------

2)Mr javed invested a total of Rs 10,000 in three different accouts , the accounts paid simple interest at an annual rate of 8%, 9%,7.5% respectively. Total interest earned for the year was Rs 845 . The amount in the 9% account was the twice the amount invested in the 7.5% account. How much did Mr.Javed invest in each account.
Equations:
Quantity Eq: E + N + S = 10000
Interest Eq: 0.08x + 0.09N + 0.075S = 845
Quatity Eq: N = 2S
-------------------------------------------
Rearrange:
E + N + S = 10000
8E+9N +7.5S=84500
0 + N - 2S = 0
----------------------
Solve this by any method you know to get:
E = 1000
N = 6000
S = 3000
============
Cheers,
Stan H.